OpenStudy (anonymous):

Find the product (X+√2) (x-√2) Show Steps

OpenStudy (watchmath):

Usint the identity \((a+b)(a-b)=a^2-b^2\) the expression can be simplify into \(x^2-(\sqrt{2})^2=x^2-2\).